ADNAN MUHAMMED

Grade 12,

Solve the equation (x^2+3x+1)(x^2+3x-3)>=5

Solve the equation (x^2+3x+1)(x^2+3x-3)>=5

Grade:

1 Answers

bharat bajaj IIT Delhi
10 years ago
Take y = x^2 + 3x
Now, (y+1)(y-3)>=5
y^2 -2y -8 >=0
(y+2)(y-4)>=0
Now y = (-inf,-2) (4,inf)

x^2 + 3x <= -2
(x+2)(x+1)<=0
x belongs from (-1,-2)
x^2 + 3x >=4
(x+4)(x-1)>=0
Hence, x= (-inf,-4) (1,inf)
These are the solutions
THanks
Bharat Bajaj
askiitians faculty
IIT Delhi